OpenStudy (anonymous):

What is the slope of the line that passes through the points (1, 1) and (–1, –5)?

OpenStudy (anonymous):

\[ slope= \frac{y_2-y_1} {x_2-x_1} \]

OpenStudy (anonymous):

Can you finish it now?

OpenStudy (anonymous):

(1, 1) and (–1, –5)? \[ slope= \frac{y_2-y_1} {x_2-x_1}= \frac{-5-1} {-1-1}=3 \]